# D-Brane Effective Lagrangian in Spacetimes with Boundaries

### v2

@@
For the boundary terms, the application of T-duality symmetry to massless open and closed string fields enables us to extend the DBI Lagrangian to include the boundary unit normal vector field $n_{\mu}$ and its first derivative.
@@ 1 Introduction
To explore the boundary world-volume Lagrangian, we examine the case where the $D_{p}$-brane is entirely along the boundary. In this setup, there exists an additional vector field $n_{a}=\partial_{a}X^{\mu}n_{\mu}$ in addition to the massless fields in the DBI Lagrangian. Our objective is to incorporate this vector field and its first derivative in the DBI Lagrangian. We will demonstrate that the T-duality of the massless open and closed string fields determines the following boundary Lagrangian:
$\partial\mathbb{L}_{p}$ $=$ $-T_{p}\,e^{-\Phi}\sqrt{-\det({\tilde{G}}_{ab}+{\tilde{B}}_{ab}+F_{ab}-2\sqrt{\alpha^{\prime}}\partial_{a}X^{\mu}\partial_{b}X^{\nu}(-K_{\mu\nu}+2n_{\mu}\partial_{\nu}\Phi))}\,.$ (6)
In proving the T-duality of the above Lagrangian, we consider a particular background for the massless closed string fields, where the circular reduction respects the $U(1)\times U(1)$ symmetry. Some parts of the above Lagrangian have already been found in [15] through an analysis of the T-duality of massless closed string fields. The inclusion of higher-derivative terms of $n_{\mu}$, such as $\nabla K$, presumably involves an infinite number of $F$ terms that may not be written in a compact form like the Lagrangian above. We are not interested in such couplings for this paper. However, the higher-order couplings that do not involve the vector $n_{\mu}$ are those that appear in the bulk world-volume Lagrangian, which we anticipate to find at order $\alpha^{\prime}$. In other words, the boundary world-volume Lagrangian includes all couplings in the bulk Lagrangian as well as other couplings involving the vector field $n_{\mu}$.
@@ 2.2 T-Duality Constraint on the Basis
Then the reduction of the DBI factor in (19) is the same as the second term in (9). The above reduction of the pull-back metric allows the covariant derivative $\nabla_{\tilde{a}}$ constructed from the pull-back metric $\tilde{G}_{{\tilde{a}}{\tilde{b}}}$, which involves the inverse of this metric in the Christoffel connection, to have an expansion in terms of $\partial_{\sigma}S\partial_{\sigma}S$. This allows the reduction of $\Omega$ and $\nabla F$ to be
@@ 3 Boundary Lagrangian
When a $D_{p}$-brane is entirely embedded in the boundary, the world-volume theory should include all the massless fields as well as the unit vector field orthogonal to the boundary. The pull-back of the massless closed string fields as well as the first derivative of the massless open string fields appear nonlinearly in the DBI action. This raises the question of whether it is possible to include the normal vector and its first derivative, which is the extrinsic curvature, into the DBI action. To find the answer to this question, one may consider T-duality on massless closed string fields or T-duality on massless open and closed string fields.
This question has been addressed in [15] by considering the massless closed string fields. It has been argued in [15] that T-duality does not allow terms such as $n\tilde{B}\nabla\tilde{B}$, $n\tilde{B}\tilde{B}\tilde{B}\nabla\tilde{B}$, and so on, in the boundary action. The terms that include the trace of the pull-back of the extrinsic curvature seem to arise from incorporating the extrinsic curvature under the square root in the DBI action and subsequently expanding it. Additionally, due to a restriction on selecting the base space background such that the circular reduction should have the symmetry $U(1)\times U(1)$ [15], the couplings in which the indices of the extrinsic curvature are contracted with $\tilde{B}$’s cannot be fixed by T-duality. By examining the T-duality of massless open and closed string fields in this section, we find that such couplings can also be derived by expanding the following extension of the DBI action:
$\partial\mathbb{L}_{p}$ $=$ $-T_{p}\,e^{-\Phi}\sqrt{-\det({\tilde{G}}_{ab}+{\tilde{B}}_{ab}+F_{ab}+2\sqrt{\alpha^{\prime}}\partial_{a}X^{\mu}\partial_{b}X^{\nu}K^{\prime}_{\mu\nu})}\,,$ (33)
where $K^{\prime}_{\mu\nu}\equiv K_{\mu\nu}-2n_{\mu}\partial_{\nu}\Phi$.
@@
Since there is no vector field in these reductions, the corresponding reduction of the Lagrangian in (33) has the symmetry $U(1)\times U(1)$. The above reduction produces the following reduction for $K^{\prime}_{\mu\nu}$:
$K^{\prime}_{\mu\nu}=\left(\matrix{\bar{K}^{\prime}_{{\tilde{\mu}}{\tilde{\nu}}}&0\!\!\!\!\!&\cr 0&0\!\!\!\!\!&}\right),$ (36)
where $\bar{K}^{\prime}_{{\tilde{\mu}}{\tilde{\nu}}}=\bar{K}_{{\tilde{\mu}}{\tilde{\nu}}}-2n_{{\tilde{\mu}}}\partial_{\tilde{\nu}}\bar{\phi}$ and $\bar{K}_{{\tilde{\mu}}{\tilde{\nu}}}=\tilde{\nabla}_{\tilde{\mu}}n_{{\tilde{\nu}}}-n_{\tilde{\mu}}n^{\tilde{\rho}}\tilde{\nabla}_{\tilde{\rho}}n_{{\tilde{\nu}}}$ is the extrinsic curvature in the base space. $\bar{K}^{\prime}_{{\tilde{\mu}}{\tilde{\nu}}}$ is invariant under T-duality.
@@
$\partial{L}_{p}^{w}=-T_{p}\,e^{-\bar{\phi}+\varphi/4}\sqrt{-\det\Big{[}\partial_{\tilde{a}}X^{\tilde{\mu}}\partial_{\tilde{b}}X^{\tilde{\nu}}(\bar{g}_{{\tilde{\mu}}{\tilde{\nu}}}+\bar{b}_{{\tilde{\mu}}{\tilde{\nu}}}+2\sqrt{\alpha^{\prime}}\bar{K}^{\prime}_{{\tilde{\mu}}{\tilde{\nu}}})+F_{{\tilde{a}}{\tilde{b}}}+e^{-\varphi}\partial_{\tilde{a}}A_{y}\partial_{\tilde{b}}A_{y}\Big{]}}\,.$ (37)
@@
$\partial{L}_{p-1}^{t}=-T_{p-1}\,e^{-\bar{\phi}-\varphi/4}\sqrt{-\det\Big{[}\partial_{\tilde{a}}X^{\tilde{\mu}}\partial_{\tilde{b}}X^{\tilde{\nu}}(\bar{g}_{{\tilde{\mu}}{\tilde{\nu}}}+\bar{b}_{{\tilde{\mu}}{\tilde{\nu}}}+2\sqrt{\alpha^{\prime}}\bar{K}^{\prime}_{{\tilde{\mu}}{\tilde{\nu}}})+F_{{\tilde{a}}{\tilde{b}}}+e^{\varphi}\partial_{\tilde{a}}S\partial_{\tilde{b}}S\Big{]}}\,.$ (38)
@@
The above T-duality symmetry does not fix the coefficient of $K^{\prime}_{\mu\nu}$ in the modified DBI action in (33). This coefficient is fixed in [15] by imposing T-duality on the combination of bulk and boundary actions when the $D_{p}$-brane extends in the bulk of spacetime and ends on the boundary. In this case, the bulk of the D-brane is T-dual up to some total derivative term in the base space. These anomalous total derivative terms then transfer to the boundary by using the Stokes theorem. Then, the boundary couplings should be such that their T-duality transformation cancels the anomalous term. In this way, the boundary coupling produces the extrinsic curvature that is in harmony with the bulk action. The coefficient of the extrinsic curvature in (33) is the one fixed in [15].
@@ 4 Discussion
$-T_{p}\sqrt{\alpha^{\prime}}\int d^{p}\tau\,e^{-\Phi}\sqrt{-\det({\tilde{G}}_{\bar{a}\bar{b}}\!+\!{\tilde{B}}_{\bar{a}\bar{b}}\!+\!F_{\bar{a}\bar{b}}\!+\!2\sqrt{\alpha^{\prime}}\tilde{K}^{\prime}_{\bar{a}\bar{b}})}[1\!+\!O_{bulk}(\alpha^{\prime})\!+\!O_{n}(\alpha^{\prime})],$
where $K^{\prime}_{\mu\nu}\equiv K_{\mu\nu}-2n_{\mu}\partial_{\nu}\Phi$. The corrections $O_{bulk}(\alpha^{\prime})$ for the bulk and boundary are the same. In the pull-back of the spacetime fields in the bulk, $X^{\mu}=X^{\mu}(\sigma)$, and in the boundary, $X^{\mu}=X^{\mu}(\tau)$. The bulk world-volume coordinates are $\sigma^{0},\cdots,\sigma^{p}$, and the boundary world-volume coordinates are $\tau^{0},\cdots,\tau^{p-1}$. Each term in the $\alpha^{\prime}$-corrections $O_{n}(\alpha^{\prime})$ should involve the unit vector $n_{\mu}$.

### v3

@@
For the boundary terms, applying T-duality symmetry to the massless open string field allows us to extend the DBI Lagrangian to incorporate the extrinsic curvature of the boundary.
@@ 1 Introduction
To explore the boundary world-volume Lagrangian, we examine the case where the $D_{p}$-brane is entirely along the boundary. In this setup, there exists an additional vector field $n_{a}=\partial_{a}X^{\mu}n_{\mu}$ in addition to the massless fields in the DBI Lagrangian. Our objective is to incorporate this vector field and its first derivative in the DBI Lagrangian. We will demonstrate that the T-duality of the massless open string fields determines the following boundary Lagrangian:
$\partial\mathbb{L}_{p}$ $=$ $-T_{p}\,e^{-\Phi}\sqrt{-\det({\tilde{G}}_{ab}+{\tilde{B}}_{ab}+F_{ab}+2\sqrt{\alpha^{\prime}}\partial_{a}X^{\mu}\partial_{b}X^{\nu}K_{\mu\nu})}\,,$ (6)
where $K_{\mu\nu}$ is the extrinsic curvature (3), in which the covariant derivative is replaced by the partial derivative. Some parts of the above Lagrangian have already been found in [15] through an analysis of the T-duality of massless closed string fields. The inclusion of higher-derivative terms of $n_{\mu}$, such as $\partial K$, presumably involves an infinite number of $F$ terms that may not be written in a compact form like the Lagrangian above. We are not interested in such couplings for this paper. However, the higher-order couplings that do not involve the vector $n_{\mu}$ are those that appear in the bulk world-volume Lagrangian, which we anticipate to find at order $\alpha^{\prime}$. In other words, the boundary world-volume Lagrangian includes all couplings in the bulk Lagrangian as well as other couplings involving the vector field $n_{\mu}$.
@@ 2.2 T-Duality Constraint on the Basis
Then the reduction of the DBI factor in (19) is the same as the second term in (9). The above reduction of the pull-back metric allows the covariant derivative $\nabla_{\tilde{a}}$ constructed from the pull-back metric $\tilde{G}_{{\tilde{a}}{\tilde{b}}}$, which involves the inverse of this metric in the Christoffel connection, to have an expansion in terms of $\partial_{\tilde{a}}S\partial_{\tilde{b}}S$. This allows the reduction of $\Omega$ and $\nabla F$ to be
@@ 3 Boundary Lagrangian
When a $D_{p}$-brane is entirely embedded in the boundary, the world-volume theory should include all the massless fields as well as the unit vector field orthogonal to the boundary. The pull-back of the massless closed string fields as well as the first derivative of the massless open string fields appear nonlinearly in the DBI action. This raises the question of whether it is possible to include the normal vector and its first derivative, which is the extrinsic curvature, into the DBI action. To find the answer to this question, one may consider T-duality on massless closed string fields, or T-duality on massless open string fields.
This question has been addressed in [15] by considering the massless closed string fields. It has been argued in [15] that T-duality does not allow terms such as $n\tilde{B}\nabla\tilde{B}$, $n\tilde{B}\tilde{B}\tilde{B}\nabla\tilde{B}$, and so on, in the boundary action. The terms that include the trace of the pull-back of the extrinsic curvature seem to arise from incorporating the extrinsic curvature under the square root in the DBI action and subsequently expanding it. Additionally, due to a restriction on selecting the base space background such that the circular reduction should have the symmetry $U(1)\times U(1)$ [15], the couplings in which the indices of the extrinsic curvature are contracted with $\tilde{B}$’s cannot be fixed by T-duality. By examining the T-duality of massless open string fields in this section, we find that such couplings can also be derived by expanding the following extension of the DBI action:
$\partial\mathbb{L}_{p}$ $=$ $-T_{p}\,e^{-\Phi}\sqrt{-\det({\tilde{G}}_{ab}+{\tilde{B}}_{ab}+F_{ab}+2\sqrt{\alpha^{\prime}}\partial_{a}X^{\mu}\partial_{b}X^{\nu}K_{\mu\nu})}\,.$ (33)
where $K_{\mu\nu}$ is given in (3), in which the covariant derivative is replaced by the partial derivative. We assume the closed string fields are constant, and the partial derivative of the normal vector to the boundary is not zero, i.e., $K_{\mu\nu}\neq 0$.
@@
Since there is no vector field in these reductions, the corresponding reduction of the Lagrangian in (33) has the symmetry $U(1)\times U(1)$. The above reduction produces the following reduction for $K_{\mu\nu}$:
$K_{\mu\nu}=\left(\matrix{\bar{K}_{{\tilde{\mu}}{\tilde{\nu}}}&0\!\!\!\!\!&\cr 0&0\!\!\!\!\!&}\right),$ (36)
where $\bar{K}_{{\tilde{\mu}}{\tilde{\nu}}}=\partial_{\tilde{\mu}}n_{{\tilde{\nu}}}-n_{\tilde{\mu}}n^{\tilde{\rho}}\partial_{\tilde{\rho}}n_{{\tilde{\nu}}}$ is the extrinsic curvature in the base space, which is invariant under T-duality.
@@
$\partial{L}_{p}^{w}=-T_{p}\,e^{-\bar{\phi}+\varphi/4}\sqrt{-\det\Big{[}\partial_{\tilde{a}}X^{\tilde{\mu}}\partial_{\tilde{b}}X^{\tilde{\nu}}(\bar{g}_{{\tilde{\mu}}{\tilde{\nu}}}+\bar{b}_{{\tilde{\mu}}{\tilde{\nu}}}+2\sqrt{\alpha^{\prime}}\bar{K}_{{\tilde{\mu}}{\tilde{\nu}}})+F_{{\tilde{a}}{\tilde{b}}}+e^{-\varphi}\partial_{\tilde{a}}A_{y}\partial_{\tilde{b}}A_{y}\Big{]}}\,.$ (37)
@@
$\partial{L}_{p-1}^{t}=-T_{p-1}\,e^{-\bar{\phi}-\varphi/4}\sqrt{-\det\Big{[}\partial_{\tilde{a}}X^{\tilde{\mu}}\partial_{\tilde{b}}X^{\tilde{\nu}}(\bar{g}_{{\tilde{\mu}}{\tilde{\nu}}}+\bar{b}_{{\tilde{\mu}}{\tilde{\nu}}}+2\sqrt{\alpha^{\prime}}\bar{K}_{{\tilde{\mu}}{\tilde{\nu}}})+F_{{\tilde{a}}{\tilde{b}}}+e^{\varphi}\partial_{\tilde{a}}S\partial_{\tilde{b}}S\Big{]}}\,.$ (38)
@@
The above T-duality symmetry does not fix the coefficient of $K_{\mu\nu}$ in the modified DBI action in (33). This coefficient is fixed in [15] by imposing T-duality on the combination of bulk and boundary actions when the $D_{p}$-brane extends in the bulk of spacetime and ends on the boundary. In this case, the bulk of the D-brane is T-dual up to some total derivative term in the base space. These anomalous total derivative terms then transfer to the boundary by using the Stokes theorem. Then, the boundary couplings should be such that their T-duality transformation cancels the anomalous term. In this way, the boundary coupling produces the extrinsic curvature that is in harmony with the bulk action. The coefficient of the extrinsic curvature in (33) is the one fixed in [15].
@@ 4 Discussion
$-T_{p}\sqrt{\alpha^{\prime}}\int d^{p}\tau\,e^{-\Phi}\sqrt{-\det({\tilde{G}}_{\bar{a}\bar{b}}\!+\!{\tilde{B}}_{\bar{a}\bar{b}}\!+\!F_{\bar{a}\bar{b}}\!+\!2\sqrt{\alpha^{\prime}}\tilde{K}_{\bar{a}\bar{b}})}[1\!+\!O_{bulk}(\alpha^{\prime})\!+\!O_{n}(\alpha^{\prime})].$
The corrections $O_{bulk}(\alpha^{\prime})$ for the bulk and boundary are the same. In the pull-back of the spacetime fields in the bulk, $X^{\mu}=X^{\mu}(\sigma)$, and in the boundary, $X^{\mu}=X^{\mu}(\tau)$. The bulk world-volume coordinates are $\sigma^{0},\cdots,\sigma^{p}$, and the boundary world-volume coordinates are $\tau^{0},\cdots,\tau^{p-1}$. Each term in the $\alpha^{\prime}$-corrections $O_{n}(\alpha^{\prime})$ should involve the unit vector $n_{\mu}$.